A one-year call option contract on Cheesy Poofs Co. stock sells for $1,170. In one year, the stock will be worth $49 or $70 per share. The exercise price on the call option is $62. What is the current value of the stock if the risk-free rate is 3 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Stock sales for =$1170

Shares per contract =100

Call value. =1170/100 =$11.70per share

Current stock value. =Delta * Call value *Amount borrowed

Delta = 70-49/70-62 = 21/8 = 2.625

Amount borrowed = 62/(1+3%) = 60.1942

Current stock value = 2.625 *11.70 +60.194

=30.7125+60.194 =$ 90.9065

Therefore the current stock value is $ 90.91